OpenStudy (df001):

Help with an Algebra II question! (Subtracting fractions with unlike denominator.)

OpenStudy (df001):

\[\frac{ S-1 }{ S } - \frac{ T+1 }{ T }\]

jimthompson5910 (jim_thompson5910):

what's the LCD in this case?

OpenStudy (df001):

I have trouble figuring out the LCD, I was never good at it. Please teach me :(

OpenStudy (df001):

I thought you multiple the fractions by the opposite denominators

jimthompson5910 (jim_thompson5910):

the denominators are S and T they have nothing in common except for 1, so we can multiply them to get the LCD the LCD is simply `ST` or `TS`

jimthompson5910 (jim_thompson5910):

to add or subtract fractions, the denominators must be the same. They aren't in this case, so we have to get them all equal to the LCD

jimthompson5910 (jim_thompson5910):

the first denominator is S we want it to be equal to ST or TS it's missing a T, so multiply top and bottom of the first fraction by T \[\Large \frac{ S-1 }{ S } - \frac{ T+1 }{ T }\] \[\Large \frac{\color{red}{T}}{\color{red}{T}}\times\frac{ S-1 }{ S } - \frac{ T+1 }{ T }\] \[\Large \frac{ \color{red}{T}(S-1) }{ \color{red}{T}S } - \frac{ T+1 }{ T }\]

OpenStudy (df001):

If in other cases like one denominator is 3a+12 and the other is a+4, how can I find the LCD in this?

jimthompson5910 (jim_thompson5910):

notice how the `T/T` is a fancy form of 1. Multiplying by 1 doesn't change the fraction

OpenStudy (df001):

As to why do we want it to equal ST or TS?

jimthompson5910 (jim_thompson5910):

you would factor 3a+12 to get 3(a+4) we really have the denominators 3(a+4) and a+4 so the LCD is 3(a+4). The unique factors are 3 and (a+4)

OpenStudy (df001):

Can I get a brief definition of what an lcd is

jimthompson5910 (jim_thompson5910):

LCD = lowest common denominator it's the LCM of the denominators

OpenStudy (df001):

Im not sure what it really means but, i thought it was the lowest number that can go into two numbers

jimthompson5910 (jim_thompson5910):

for example 1/2 + 2/3 the LCD is 6 since the LCM of 2 and 3 is 6

jimthompson5910 (jim_thompson5910):

6 is the lowest multiple both 2 and 3 have in common

OpenStudy (df001):

Did you add the denominators together in 1/2+2/3

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

list the multiples of 2 2, 4, 6, 8, 10, 12, ... list the multiples of 3 3, 6, 9, 12, 15, 18, ... the common multiples are 6, 12, 18, ... the number 6 is the smallest multiple

jimthompson5910 (jim_thompson5910):

so 6 is the LCM (lowest common multiple)

OpenStudy (df001):

Back to the image 5 minutes ago, must you look at the first fraction first to know what to times the numerator and denominator by?

jimthompson5910 (jim_thompson5910):

you mean back with the one with S and T in it?

OpenStudy (df001):

I see the first fraction is times by T thus ST

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (df001):

But, if I do that to the first fraction, should I do that to the second fraction ?

OpenStudy (df001):

mulitple T by S

jimthompson5910 (jim_thompson5910):

you'll do something similar, but with S instead

OpenStudy (df001):

thus, they have the same denominator ST

jimthompson5910 (jim_thompson5910):

you'll multiply top and bottom of the second fraction by S

jimthompson5910 (jim_thompson5910):

yes that's the ultimate goal: to get the denominators equal to the LCD (so they are all the same)

OpenStudy (df001):

Oh, The answer I got for the ST question is -1T+1S/ST

OpenStudy (df001):

If I am correct about the numerator

jimthompson5910 (jim_thompson5910):

so we have this so far after getting each denominator equal to the LCD \[\Large \frac{ S-1 }{ S } - \frac{ T+1 }{ T }\] \[\Large \frac{\color{red}{T}}{\color{red}{T}}\times\frac{ S-1 }{ S } - \frac{ T+1 }{ T }\] \[\Large \frac{ \color{red}{T}(S-1) }{ \color{red}{T}S } - \frac{ T+1 }{ T }\] \[\Large \frac{ T(S-1) }{ TS } - \frac{\color{red}{S}}{\color{red}{S}}\times\frac{ T+1 }{ T }\] \[\Large \frac{ T(S-1) }{ TS } - \frac{ \color{red}{S}(T+1) }{ \color{red}{S}T }\] \[\Large \frac{ T(S-1) }{ TS } - \frac{ S(T+1) }{ TS }\] agreed?

OpenStudy (df001):

Yes but, I mines is T(S-1)-S(T+1)/TS

OpenStudy (df001):

im not sure if thats the same as the last step

jimthompson5910 (jim_thompson5910):

yeah you'll have T(S-1)-S(T+1) all over TS as one of your steps

OpenStudy (df001):

I combined the denominator

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ T(S-1) }{ TS } - \frac{ S(T+1) }{ TS }\] turns into \[\Large \frac{ T(S-1) - S(T+1) }{ TS }\]

OpenStudy (df001):

If I multiply t and s into what's in the parenthesis do I get \[\frac{ -1t+1s }{ st }\]

jimthompson5910 (jim_thompson5910):

good, which is equivalent to \[\Large \frac{S-T}{ST}\]

jimthompson5910 (jim_thompson5910):

the TS terms up in the numerator will cancel (since 1TS - 1TS = 0TS = 0)

OpenStudy (df001):

Why is it S-T

jimthompson5910 (jim_thompson5910):

oh wait, that S should be negative

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ T(S-1) - S(T+1) }{ TS }\] \[\Large \frac{ TS-T - ST-S }{ TS }\] \[\Large \frac{ -T -S }{ TS }\]

jimthompson5910 (jim_thompson5910):

hopefully all that makes sense?

OpenStudy (df001):

Oh, I see now

OpenStudy (df001):

I got confused by the -S

OpenStudy (df001):

thank you man

jimthompson5910 (jim_thompson5910):

sure thing